The lawn in front of Pythagoras Jr. High is in the shape of a rectangle 24 m long and 10 m wide. How many meters shorter is your walk if you walk diagonally across the lawn rather than along two sides of it?
Answer by penguin18(7)

Use pythagorean theorem: a^2+b^2=c^2
plug in the two known values for a and b: 24^2+10^2=c^2
square 24 and 10: 576+100=c^2
Add them together: 676=c^2
Take the square root of 676: c=26
Subtract 26 from the total distance you would have to travel to get your answer:
34-26=8 meters shorter.